Newton's law of cooling states that the temperature of an object changes at a rate proportional to the difference between its temperature and that of its surroundings. Suppose that the temperature of a cup of coffee obeys Newton's law of cooling. If the coffee has a temperature of 195 degrees Fahrenheit when freshly poured, and 1 minutes later has cooled to 179 degrees in a room at 80 degrees, determine when the coffee reaches a temperature of 139 degrees.
The coffee will reach a temperature of 139 degrees in ____________ minutes.
Answer #1

As per newton's Law of cooling

\(\frac{d T}{d t}=-k(T-S)\)

where \(T=\) current temperature and

\(S=\) ambient temperature

\(\Rightarrow \frac{d T}{T-S}=-k d t\)

Solve the differential equation gives,

\(\ln (T-S)=-k t+C\)

\(\Rightarrow T-S=e^{-k t+C}\)

\(\Rightarrow T(t)=S+e^{-k t+C}\)

\(\Rightarrow T(t)=S+\left(T_{0}-S\right) e^{-k t}\)

where \(T_{0}=\) initial temperature at \(t=0\) \(\Rightarrow T(t)=80+(195-80) e^{-k t}\)

$$ T(t)=80+115 e^{-k t} $$

\(\Rightarrow 179=80+115 e^{-k k}\)

\(99=115 e^{-k}\)

\(e^{-k}=\frac{99}{115}\)

\(-k=\ln \left(\frac{99}{115}\right)\)

\(k=-\ln \left(\frac{99}{115}\right)\)

\(k \approx 0.1498\)

Therefore, \(T(t)=80+115 e^{-0.1498 t}\)

\(139=80+115 e^{-0.1498 t}\)

\(59=115 e^{-0.1498 t}\)

\(e^{-0.1498 t}=\frac{59}{115}\)

\(-0.1498 t=\ln \left(\frac{59}{115}\right)\)

\(t=-\frac{1}{0.1498} \ln \left(\frac{59}{115}\right)\)

\(\approx 4.455\) minutes